Jill is 22 kilometers away from Joe. Both begin to walk toward each other at the same time. Jill walks at 3km/hr. They meet in 4 hours. How fast is Joe walking?
Answer by mananth(16949)

Jill 3km/h
Joe xkm/h
towrads each other so combined speed = x+3 km/h
distance = 22 km
t= distance/speed
time = 4 h
22/(x+3)=4
22=4(x+3)
22=4x+12
4x=10
x=2.5 km/h Joe's speed